what are the odds against drawing a red face card from a standard 52-card deck?
Answer by checkley75(3666)

thre are 4 face cards in hearts & diamonds thus there are 8 out of 52 red face cards or
8/52=4/13=.1538461 probability of drawing a red face card. (15.38461%)